OpenStudy (babybritt002):

need help step by step help needed 1/r-5=7/2r

Vocaloid (vocaloid):

does your question look like this?

OpenStudy (babybritt002):

yes @Vocaloid

Vocaloid (vocaloid):

start by cross multiplying

OpenStudy (anthonyym):

Use cross multiplication. 1*2r = 7(r-5)

Vocaloid (vocaloid):

^ know where to go from there?

OpenStudy (babybritt002):

no @Vocaloid

Vocaloid (vocaloid):

distribute the 7. 7(r-5) = ?

OpenStudy (babybritt002):

so 7r+2

OpenStudy (babybritt002):

@Vocaloid

jhonyy9 (jhonyy9):

no multiplie 7 by -5 too

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

how distribution works: A(B-C) = A*B - A*C so, following that same pattern, what is 7(r-5) = ?

OpenStudy (babybritt002):

7r-35

jhonyy9 (jhonyy9):

yes

Vocaloid (vocaloid):

good! now we have 2r = 7r - 35 subtract 2r from each side. what do you get?

OpenStudy (babybritt002):

so 5r-35

Vocaloid (vocaloid):

almost, don't forget the 0 on the left side 0 = 5r - 35 now add 35 to both sides

OpenStudy (jeansweets):

r=7

OpenStudy (babybritt002):

35=5r

Vocaloid (vocaloid):

good, now divide both sides by 5

OpenStudy (jeansweets):

you have to solve the rational equation by combining expressions and isolating the varible r

OpenStudy (babybritt002):

ok so 35/5 = 7

OpenStudy (babybritt002):

so r=7

OpenStudy (babybritt002):

@Vocaloid

Vocaloid (vocaloid):

good

OpenStudy (jeansweets):

yes